RC meeting and Run Plan, Sunday March 12th
Return to the RC Meeting Calendar
Contents
RC Report
- RC: Holly
Summary
- RC Daily report: https://logbooks.jlab.org/entry/4148180
We established a stable running current around 35uA yesterday morning. Just before noon, we lost beam due to problems with 1L04. Beam returned at 35uA around 1:30pm and continued until 3pm when the beam was taken away for orbit checks due to an increase in trips. No obvious result was found and beam resumed after the start of swing. The target group worked to improve the polarization during the day shift yesterday. Some analysis was done to study the convection settings (https://logbooks.jlab.org/entry/4147689). A laser was added to the system to increase power to 180W on the cell (https://logbooks.jlab.org/entry/4147707). Just before the start of swing, Hall B requested the IHWP change to IN (https://logbooks.jlab.org/entry/4147754). Around 6pm, there was a problem with the DAQ due to configuration errors with the MPDs in the SBS GEMs. Experts jumped in quickly to power cycle crates, but the root cause was not quickly resolved nor understood so we commenced running without the SBS GEMs in the DAQ. We are looking for another window without beam when we can troubleshoot the crate issue. It was noted earlier in the day that the target RTD was noisy so Bill changed the algorithm for picking up the max temperature in the feedback loop (https://logbooks.jlab.org/entry/4147972). Beam continued smoothly throughout the remainder of swing and the shortened owl shift.
Day summary: https://logbooks.jlab.org/entry/4148155 Owl summary: https://logbooks.jlab.org/entry/4148002 Swing summary: https://logbooks.jlab.org/entry/4147865
Recent Issues
- While we request and are able to run at 45uA, accelerator has us running at 35uA for stability with Hall C online.
- Restarting the target ladder control window: https://logbooks.jlab.org/entry/4147909
- Paul King adjust the PITA offset with IHWP IN: https://logbooks.jlab.org/entry/4147914, https://logbooks.jlab.org/entry/4147935
Current Status
- Latest carbon hole image: https://logbooks.jlab.org/entry/4147220
- Running 45uA on pol-3He
Run Plan
- Production on 3He!
- GENII-3Stream config (currently no SBSGEMs until we can put them back in)
- PS1=2, PS3=0, PS5=6
- Useful guide from Bill on rates to compare when bringing beam back: https://logbooks.jlab.org/entry/4144886
Data Collection and Polarization Status
- Currently 41%: https://logbooks.jlab.org/entry/4148138
- Data status at 4th-pass: 49.3 C (uncorrected)